OpenStudy (anonymous):

x^3-x^2-6x>/0 inequality

OpenStudy (turingtest):

\[x^3-x^2-6x \ge 0\]\[x(x^2-x-6) \ge0\]\[x(x-3)(x+2) \ge 0\]the solutions are\[x \ge 0\]\[x \ge 3\]\[x \ge -2\] since the second expression implies the others it is our limiting agent here, so\[\left\{x: x \ge 3 \right\}\]